Target (1/31/22) Wal-Mart (1/31/22) Income Statement Data for Year Net sales $67,000 $410,000 Cost of goods...

Target
(1/31/22)

Wal-Mart
(1/31/22)

Income Statement Data for Year

Net sales

$67,000

$410,000

Cost of goods sold

44,890

287,000

Selling and administrative expenses

16,080

77,900

Interest expense

670

3,280

Other income

20

4,000

Income taxes

1,883

16,037

Net income

$ 3,497

$ 29,783

Target

Wal-Mart

Balance Sheet Data (End of Year)

Current assets

$21,160

$ 45,360

Noncurrent assets

27,000

111,000

Total assets

$48,160

$156,360

Current liabilities

$11,500

$ 56,000

Long-term liabilities

22,212

37,816

Total stockholders’ equity

14,448

62,544

Total liabilities and stockholders’ equity

$48,160

$156,360

Net cash provided by operating activities

$5,200 $23,400

Cash paid for capital expenditures

$3,700 $11,800

Dividends declared and paid on common stock

$450 $3,400

Weighted-average common shares outstanding (millions)

650 3,950



For each company, compute these values and ratios. (All dollars are in millions.) (Round Current ratio and Earnings per share to 2 decimal places, e.g. 15.25 and Debt to assets ratio to 1 decimal place, e.g. 78.9%. If answer is negative enter it with a negative sign preceding the number e.g. -15,000 or in parentheses e.g. (15,000).)

Target Wal-Mart
(a)

Working capital

$ enter a dollar amount for working capital for Target $ enter a dollar amount for working capital for Wal-Mart
(b)

Current ratio

enter the current ratio for Target rounded to 2 decimal places :1 enter the current ratio for Wal-Mart rounded to 2 decimal places :1
(c)

Debt to assets ratio

enter debt for assets ratio for Target in percentages rounded to 1 decimal place % enter debt for assets ratio for Wal-Mart in percentages rounded to 1 decimal place %
(d)

Free cash flow

$ enter a dollar amount for free cash flow for Target $ enter a dollar amount for free cash flow for Wal-Mart
(e)

Earnings per share

$ enter a dollar amount for earnings per share for Target rounded to 2 decimal place $ enter a dollar amount for earnings per share for Wal-Mart rounded to 2 decimal place
(f)

Which company has better liquidity?

select between Target or Wal-Mart                                                          TargetWal-Mart

Which company has better solvency?

select between Target or Wal-Mart                                                          TargetWal-Mart
Answer #1

Calculate following

Target Wal-mart
1. working capital 21160-11500 = 9660 45360-56000 = -10640
2. Current ratio 21160/11500 = 1.84:1 45360/56000 = 0.81:1
3. Debt to assets ratio 33712/48160 = 70% 93816/156360 = 60%
4. Free cash flow 1050 8200
5. Earning per share 5.38 7.54
(f)

Which company has better liquidity?

Target

Which company has better solvency?

Wal-mart